OpenStudy (anonymous):

Square root of 13 over 11 .

OpenStudy (anonymous):

\[\sqrt{13}/11\]

OpenStudy (vijay):

\[\sqrt{13/11}\]

OpenStudy (anonymous):

1.087

OpenStudy (anonymous):

or do you mean (the square root of 13) divided by 11?

OpenStudy (anonymous):

- Yes , that .

OpenStudy (anonymous):

simplified or the number value?

OpenStudy (anonymous):

Both

jimthompson5910 (jim_thompson5910):

\[\sqrt{\frac{13}{11}}\] \[\frac{\sqrt{13}}{\sqrt{11}}\] \[\frac{\sqrt{13}*\sqrt{11}}{\sqrt{11}*\sqrt{11}}\] \[\frac{\sqrt{13*11}}{\sqrt{11*11}}\] \[\frac{\sqrt{143}}{\sqrt{121}}\] \[\frac{\sqrt{143}}{11}\] \[\frac{\sqrt{143}}{11}\approx 1.087\]

OpenStudy (anonymous):

simplified is \[\sqrt{13/11}\] value=.327777

OpenStudy (mathsolver):

0.327777 (to 6dp) or \[\sqrt{13}/11\] as a surd

OpenStudy (mathsolver):

if you want more digits they are 0.327777388678544481192656478860954176931936052167749655700950277838833358935728222291329007456226428879410380184005